OpenStudy (anonymous):

Calculate the square roots of the complex number 5 - 12i. ps.:the result is 3 - 2i; -3 + 2i but how can I do it?

OpenStudy (zarkon):

\[(a+bi)^2=5-12i\] \[a^2-b^2+2abi=5-12i\] solve \[a^2-b^2=5\] and \[2ab=12\] for a,b

OpenStudy (zarkon):

2ab=-12